The following data were drawn from the records of Perez Corporation. Planned volume for year (static budget) Standard directComplete this question by entering your answers in the tabs below. Req A Req B Reqc Req D Req E to G Prepare a materials variComplete this question by entering your answers in the tabs below. Req A Req B Reqc Req D Req E to G Calculate the materialsComplete this question by entering your answers in the tabs below. Req A Req B Reqc Req D Req E to G Prepare a labor varianceComplete this question by entering your answers in the tabs below. Req A Req B Reqc Req D Req E to G Calculate the labor pricComplete this question by entering your answers in the tabs below. Req A Req B Reqc Req D Req E to G Calculate the predetermi
Answer #1

A.

Material Variance Information Table
Standard Price $       1.50 per pound
Actual Price $       2.00 per pound
Standard Quantity for flexible budget 13650 pounds =3900*3.5
Actual Quantity used 12090 pounds =3900*3.1

B.
Material price Variance = (Actual Price - Standard Price) x Actual Quantity
= ($2 - $1.50) x 12090 = $6045 (U)

Material Quantity Variance = (Actual Quantity - Standard Quantity) x Standard Price
= (12090 - 13650) x $1.50 = $2340 (F)

C.

Labor Variance Information Table
Standard Price $       3.30 per hour
Actual Price $       2.70 per hour
Standard hours for flexible budget 8970 hours =3900*2.3
Actual hours used 10140 hours =3900*2.6

D.
Labor Rate Variance = (Actual price - Standard Price) x Actual hours
= ($2.70 - $3.30) x 10140 = $6084 (F)

Labor Efficiency Variance = (Actual hours - Standard hours) x Standard Price
= (10140 - 8970) x $3.30 = $3861 (U)

E. Predetermined Overhead rate = Budgeted Fixed Overhead / Planned Units
= $14620 / 3400 = $4.30 per unit

F. Fixed Cost spending Variance = Actual Fixed Cost - Standard Fixed Cost
= $10220 - $14620 = $4400 (F)

G. Fixed Cost Volume Variance = (Actual Quantity - Budgeted Quantity) x Predetermined Overhead rate
= (3900 - 3400) x $4.30 = $2150 (F)